Paromomycin sulfate is used to eradicate intestinal amoebas. A
typical dose consists of 35 mg/day for each kilogram of body mass.
The treatment can last up to 10 days. For a patient who weighs 198
pounds, how many 250-mg capsules of paromomycin sulfate are
required for a 10-day regimen?
Answer #1
  • A typical dose of Paromomycin sulfate consists of 35 mg/day for each kilogram of body mass.
  • The treatment can last up to 10 days.
  • Patient weight = 198 pounds = 89.8 kg
  • Number of 250-mg capsules of paromomycin sulfate required for a 10-day regimen = 125 capsules

Calculation:

35 mg / day for 1 kg

89.8 * 35 = 3143 mg / day for 89.8 kg

number of 250-mg capsules per day = 3143/250 = 12.5 = 12 and a half capsule

number of 250-mg capsules for 10 days = 12.5 * 10 = 125